OpenStudy (anonymous):

whats the quadratic formula for x2-3x+1=0

OpenStudy (anonymous):

do u want the solutions?

OpenStudy (anonymous):

iyess

OpenStudy (anonymous):

x=1/2(3+sqrt5) :P

OpenStudy (anonymous):

noo not right

OpenStudy (anonymous):

and x=1/2(3-sqrt5)

OpenStudy (anonymous):

it is :P

OpenStudy (anonymous):

i about to draw the answer choices

OpenStudy (anonymous):

the answer should be x= 3+/-sqrt13/2

OpenStudy (anonymous):

If ax^2 +bx +c =0 is the quadratic equation then the solution for x is x={-b±√(b^2 -4ac)} /(2a) Now for the given equation a=1, b=-3, c=-1 so x={-(-3)±√((-3)^2 -4(1)(-1))} /(2*1) ..={+3±√(9 +4))} /2 ..={3±√13}/2 So the solution set is {3+√13}/2, 3-√13}/2}

OpenStudy (anonymous):